OpenStudy (anonymous):

Choose the correct simplification of 2a0. 2a 0 2 1

OpenStudy (anonymous):

Is this one 2a?

OpenStudy (anonymous):

everything*0 is 0

OpenStudy (anonymous):

anything raised to zero is 1 so it becomes 2(a^0) = 2(1)

OpenStudy (anonymous):

ohh danggit......thx

OpenStudy (anonymous):

sooo mathwhiz99 is mukushia right?

OpenStudy (anonymous):

no. mukushia assumed your question meant 2a*0..i think you meant \(2a^0\) right?

OpenStudy (anonymous):

Remember: \[x^0 = 1\]

OpenStudy (diyadiya):

No! Anything raised to zero is 1 @mathwhiz99 is Right!! :)

OpenStudy (anonymous):

@Diyadiya , what is 0 raised to the power 0???

OpenStudy (anonymous):

0 raised to 0 is the end of the world

OpenStudy (diyadiya):

I guess thats 1 as well :S

OpenStudy (diyadiya):

Not sure though

OpenStudy (anonymous):

That is undefined.. Then how can you say anything raised to the power is 0???

OpenStudy (diyadiya):

Oh ok sorry about it then! my bad

OpenStudy (diyadiya):

there are always exceptions

OpenStudy (anonymous):

Not it is not bad it is good that you have gained something.... Ha ha ha...

OpenStudy (anonymous):

And remember one thing math is not an exception...

OpenStudy (anonymous):

lol so the answer is 1 right?

OpenStudy (diyadiya):

Well, I must say : Anything raised to the power Zero EXCEPT Zero is " 1 '' ^_^

OpenStudy (diyadiya):

Nope the answer is not 1 <<

OpenStudy (diyadiya):

\[2 a^0 = 2(1) = 2 \]

OpenStudy (anonymous):

Answer is other than A...

OpenStudy (anonymous):

So I was wrong...... the answer is 2?

OpenStudy (diyadiya):

Yes^

OpenStudy (anonymous):

yes. 2.

OpenStudy (anonymous):

Ya answer is 2...

OpenStudy (anonymous):

oh okay lol..because Iwas so confused with all these different people answering..But thank You all of you! ♥